A bag contains 6 cherry, 3 orange, and 2 lemon candies. You reach in and take 3 pieces of candy at random. Find the probability of all 3 orange. Round to 4 decimal places. Please solve and show work. Thank you very much.
Answer by stanbon(75887)

A bag contains 6 cherry, 3 orange, and 2 lemon candies. You reach in and take 3 pieces of candy at random. Find the probability of all 3 orange. Round to 4 decimal places. Please solve and show work. Thank you very much.
---------------
P(3 oranges) = 3C3/11C3 = 1/165 = 0.0061
--------------
Cheers,
Stan H.
-------------